40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 13 19:02:56 2016, Teng, Update, General, ITMX sensor 47.png34.png
    Reply  Wed Sep 14 09:33:54 2016, Steve, Update, General, ITMX magnets are stucked ITMXmagStuck.png
       Reply  Wed Sep 14 20:05:32 2016, Lydia, Update, General, ITMX magnets are stucked 
Message ID: 12494     Entry time: Wed Sep 14 20:05:32 2016     In reply to: 12491
Author: Lydia 
Type: Update 
Category: General 
Subject: ITMX magnets are stucked 

When I restarted c1susaux yesterday, I didn't know that I needed to disable the coil outputs first. So when it came back online, it attempted to damp all the vertex area optics and ITMX got stuck frown

We should make a note in the Computer Restart Procedures wiki page indicating the importance of disabling the coils before rebooting c1susaux, c1auxex, and c1auxey. Today c1auxey was rebooted properly without incident. If the slider values etc go back to their previous values on their own, is it necessary to do a BURT restore? I tried doing one for c1susaux today and there were some errors for ASC channels, but the alignment sliders went right back to the proper place after reboot yesterday.

Quote:

I believe that the UR and LR magnets are stuck. There was no earth quake at 16:18 yesterday. Something had to kick it into this position. See 4days plot

Please advise  freeing details

 

ELOG V3.1.3-